OpenStudy (anonymous):

is the answer 50 or -50? Factor: x^2 - 100

OpenStudy (anonymous):

neither

OpenStudy (anonymous):

100=10^2

OpenStudy (anonymous):

um what? can you explain the steps you used?

OpenStudy (anonymous):

x^2-100=x^2-10^2=(x-10)(x+10) difference of two squares

OpenStudy (anonymous):

the answer is (x-10)(x+10) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw